OpenStudy (anonymous):

how to solve this radical equation 3 squared 3x = 9 ?

OpenStudy (anonymous):

like this?: \[3\sqrt{3x}=9\]

OpenStudy (anonymous):

or (3^2)*3x=9

OpenStudy (anonymous):

or something else

OpenStudy (anonymous):

since you said radical equation my geuss is the first one

OpenStudy (anonymous):

so we divide by 3:square both sides:divide by three again: \[\sqrt{3x}=3\rightarrow3x=9 \rightarrow x=3\]

OpenStudy (anonymous):

yesss the first one

OpenStudy (anonymous):

thank you